North Eastern Locomotive Preservation Group

  • Increase font size
  • Default font size
  • Decrease font size
Q6 Blog

A record of events, activities and our work with the Q6

Saturday 20 February 2016

Andy reports another good day, with progress made thanks to Bryan and the three JVs, Terry, Peter and Ian.

The pistons had arrived back so the piston ring grooves were cleaned out and measured so some new rings can be made.

The junior volunteers have removed some more wet sand from the left sand box after putting some of their training into practice, making a scoop to reach the sand.

Some of the lubricator pipes went back onto the loco and the right front axle box new pipes were fitted. We will only be able to do the first two axles at the moment as others will require the wheel sets to be removed, to get access to the top of the boxes.

The cylinder cover copper rings have been annealed.

The side rod bush clearances have now all been checked and recorded and the split pins renewed.

Terry has checked all the lubrication trimmings and oil pots. This also included removing any water and contaminated oil.

The safety valves will be taken to Hopetown on Monday evening for a start to be made on overhauling them, with Neal’s assistance, .

The arch formers have been put out ready for collection to take to Carnforth for the K1 to get its new arch.

Ian Pearson is going to look after the Wednesday work party with the following jobs:

  • Fit left leading axle box new pipes, use right as an example.
  • Continue with all cab valve maintenance.
  • Continue cleaning loco.
  • Continue with sand box and re-connect any linkages/pipes as required.
  • Clean round steam reverser and nip up all glands, fill locking cylinder with oil.
  • Horn stay bolt holes (marked ) to ream out and machine new bolt to fit hole.

Wednesday 17 February 2016

Another wet and cold day, as it poured down from start to finish, but thankfully not the expected snow. Because of the bad weather forecast for the day, the JVs postponed their attendance to Saturday, but Derek, Jon, Mike Oliver and myself, led by Ian, spent another happy day with the Q6. Fortunately the pump was behaving itself as the water was coming in as fast as it was getting it out.

The piston valves were completed, with the covers left off, but with three badly worn studs still needing attention on the left hand valve housing. The horn stay bolts were checked and the two leading bolts on the left hand leading horn stay were marked up as needing attention. All the others appeared sound and rang true when struck with the hammer.

Derek finished cleaning and oiling the rods, and Mike continued with his work on the cab gauges - thanks for the packing for them (and for the valves) which enabled progress to be made.

A start was made on the sanders. The left hand cab sandbox was emptied of as much sand as our arms would reach with a plastic tub as a scoop: if someone could bring a long handled ladle on Saturday that would help clear the bottom five inches or so. The deeper into the sandbox, the wetter the sand, but we resisted using it to make a model of the Q6!

Instead, it is in a large plastic bag next to the workshop wall. The rear left hand sand pipe was removed in the hope that we would be able to empty the sand box from below, but the flap which moves to allow the sand to drop through was firmly stuck in the closed position.

Disconnecting the rodding mechanism allowed the operation of both sandboxes on the middle of the running plate to be checked, and they were working fine. So attention needs to focus on the left hand cab sandbox and freeing up the handle, rodding and flap, which may mean dismantling the whole mechanism from below. The hole in the frame through which part of the mechanism protrudes is also fouling the movement of the external lever and may need angle grinding to trim some metal off to allow the lever to move freely.


Saturday 13 February 2016

Only Ian Pearson and James Pearcy at Deviation Shed on Saturday, but a good day, albeit very cold and wet. The pump in the pit was working well with the mesh filtering out the remaining dirt from the pit and enabling the water to get away.

The slide bars arrived about 9.40 and were off loaded by Barnie, the duty shed fitter, and put into Deviation shed using the tele handler. Many thanks to him for his help. The L/H top and bottom slide bars were then fitted to the locomotive using the hydraulic lifting table. This great piece of kit meant the job was completed by lunchtime. After lunch, James measured up the 8 spots on the slide bars required by Neal and will pass this information on to him. A start was then made on the piston valves. The R/H piston valve was released from its grip on the valve rod, which had caused problems on Wednesday, and the L/H valve was then successfully fitted using strops and the pull lift. The L/H valve was repacked with the follower still to fit. The R/H valve is still to repack, for which some more 16mm packing might be required. Andy is going to speak to Neal about this on Monday night, as well as the separate packing requested by Mike Oliver for the cab gauges from last Wednesday's working party.

So some significant progress.

Jobs for Wednesday:

  • Finish off the sliders for the mesh in the pit to protect the pump from rubbish.
  • Check sanders work. The front ones might, but the back ones definitely do not. Will have to empty the sand boxes and investigate/repair.
  • Finish off piston valves but leave off the front covers.
  • Side rod bush clearances to check with feeler gauges and record sizes. Will have to move loco to get right leading and intermediate (check the cable for the lights is not trapped under loco!).
  • Remove the pipes from leading axle, from oil box to axle box, mark up & put into storage with the new flexible pipes(should be three and one with Neal to get some fittings made). There will be three ferrules in the top of the axle boxes to remove (right inner one already removed as sample to get fittings made). These must not be lost in case we have to revert back to the original set up. Look in top of axle boxes and remove any water from side cavities.
  • Check all horn stay bolts: left leading and intermediate reported as needing attention by the MPD. Mark up with yellow paint any that are loose. New bolts are on the bench behind the grinder, holes will require reaming.

Wednesday 10 February 2016

A day of steady if unspectacular progress. Joined by Terry Newman, who prepared the gaskets for the safety valves and whistle, the working party managed to insert the right hand valve, but the packing has yet to be fitted to the valve spindle. During the cleaning of the valve ports and the studs, three studs on the left hand side valve housing were marked up with badly worn threads which will need attention. Mike Oliver continued his sterling work on the cab gauges, and Derek cleaned the left hand side rods. The refurbished baffle plate was also delivered and is now in the workshop, whilst the slide bars had been collected during the week and taken away to Darlington for machining.


Saturday 6 February 2016

Andy reports that there was a really good turn out on Saturday, with eight volunteers and four JVs, so lots of work was achieved. The pistons, with the rods attached went to Sheffield last Tuesday to get the rods ground true. As no-one from NELPG was present, these were loaded by the MPD staff, so many thanks are due to all those who were involved. Hopefully the pistons will be back in the next two weeks, and then M-Machine can make some new piston rings and packings for the glands.

Peter and Trevor were once again cleaning and getting items ready for some paint. Once the cleaning of the steam heat pipe was completed, Peter painted it under the loco and tender before moving on to the steam chest casting and cylinder covers. Trevor helped get the whistle off and cut material for the JVs to use during their training session with Chris Parrish.

Neal and James did a survey of the left cross head and slide bars, in the light of comments on previous experience from Richard Pearson. After some deliberation, it was decided to remove them and ask M-Machine if they could machine the slide bars and put some white metal onto the edges of the slippers. Hopefully, we will have them back for next Saturday to start rebuilding that cylinder assembly.

Ian has done a marvellous job getting the boiler ready for the insurance exam: he has also checked that enough door joint gaskets are in stock to box the boiler up when the time comes.

Bryan and Steve were looking after our JVs. They had a really interesting day with instruction from Chris Parrish and were also having a go at painting, making a mesh to stop rubbish getting into the pump's sump, checking the side rod bush clearances using feeler gauges, and fitting the slacker pipe and valve in the cab.

Hopefully we have now come to the point with the maintenance where we are starting to turn the corner and putting things back together. While the locomotive certainly looks a lot better now we have some newly painted items, there is still a way to go yet.

Jobs for Wednesday:-

  • Continue with cab valves including whistle.
  • Make new gaskets for safety valves and whistle.
  • Clean valve ports of carbon and wire brush studs.
  • Fit piston valves and covers.
  • Fit packing to valve spindles.

Wednesday 3 February 2016

A much better day today than last Wednesday, with the sun shining, ice on the ground, the workshop heater going full blast and the pump working without a hitch. Ian, Mike, Derek and myself were joined by Les who had the NELPG car to collect some items to take back to Hopetown.

Ian continued with the boiler clean, and expects to get that finished on Saturday. Mike was working on the cab fittings and repacked and refitted the blower shut off valve on the manifold and the blower valve itself.

Barry Neesam kindly skimmed some of the injector steam and shut off valves, as well as machining two guides for the clack box and steam valve. The final two cylinder drain cocks were removed, cleaned and put with the rest of the assembly on the workshop bench. The right cylinder and ports were cleaned of carbon, although the very end face of the cylinder could do with some more of the carbon chipping off if someone can reach it. Les cleaned both cylinder covers, although the right hand one, which was more heavily carboned, still has the copper ring in situ.That needs very carefully removing and then cleaning underneath to finish that job. Ian says both copper rings will need annealing on Saturday. Finally, Derek cleaned the steam heat pipe under the locomotive and tender. There is still a small piece to complete under the tender and another short section where the pipe goes into the cab but otherwise it is ready for painting.

No cleaning of the valve cover outer faces was done, nor was the regulator handle removed from the square or the stuffing box and packing stripped. All jobs for Saturday?


Saturday 30 January 2016

Andy reports a lovely day at the Railway today, everyone came away black after cleaning carbon etc. from various bits of engine. Many thanks to Trevor, Peter, Ian & Neal for a really good day's work with lots of prep work completed ready for putting loco back together.

Some insulation for the cylinders has been obtained to be installed after the high temperature paint has been applied and the lubrication pipes been put back. Trevor has done a fine job on the piston heads. They were extremely thick in carbon but are now ready to go away to have the rods ground and new packings will be made to suit the new size of rod. Peter has continued with the lubrication pipes, cleaned the underside of the steam chest and started on the steam heat pipe under the loco.

Neal came with his own additional work list to make preparations for future work on the locomotive. This included finding the new packings for the piston rods, finding some gasket material for the safety valves and whistle etc. etc.

Ian has continued getting the boiler clean. Today he concentrated on finishing the doors and the area they sit inside the boiler, a very laborious job but a most important one if the doors are going to seal. Andy cleaned the carbon from the left cylinder and ports, then set up a line on the centre to check the slide bars and crosshead. The slide bars had a straight edge put against the side of the outer face: they are not parallel in the vertical plane, but the bottom one is good to the line. He is now having to think how to sort this out and take the wear out of the cross heads which are in three pieces. He also filled the boot of his car with the cross head splitting gear which will go to Hopetown on Monday.

The Wednesday job list is now:

  • Continue with boiler clean;
  • Continue with back head valves;
  • Clean right cylinder and ports of carbon;
  • Remove the last two cylinder drain cocks, clean and put all cocks and rods into safe storage;
  • Clean both cylinder covers, taking care not to damage copper ring on mating face;
  • Clean both valve cover outer faces;
  • Finish cleaning steam heat pipe under loco and tender (front of loco started), ready for painting on Saturday; and
  • Remove regulator handle from square, match mark before removal, also strip stuffing box & packing.

Wednesday 27 January 2016

Not the best of days yesterday. With the recent rain, the pit was quite full of water and the morning was spent trying to keep the pump operating and clearing the sludge and stones which kept blocking it up. With it also raining outside, the water was coming in as fast as we could get it pumped out! However, with Ian Pearson and Jon Bradley's skill, persistence, and willingness to plunge their hands into the murky water (not even the frog that we found was too keen on staying in it) the pit was eventually sufficiently empty to enable Derek to sweep it out, clear the central channel, and remove a considerable quantity of filth into a barrow for disposal. Hopefully though, that will make it easier in future to empty out the pit and enable work on the Q6 to be got on with rather earlier in the day than we managed to achieve.

So far as the engine itself was concerned, Mike Oliver continued his work on the cab valve maintenance, and the outside front and rear, and the inside front, cylinder drain cocks were taken off, cleaned and checked - they are now on the workshop bench in the appropriate places next to the rods which Les had previously removed. The inside rear ones await removal and cleaning. In between pump clearing activity, Jon managed to get some more work done on the mudhole doors, and these are now estimated at 99% complete. They are on the running plate.

No work was done however, on the slide bars or on cleaning the frames and wheels.

There is therefore plenty to be getting on with for this Saturday's working party, which Andy Lowes will be leading. Any one planning to attend of whom he is not already aware, please let him know. The NELPG car will not be available as it will be at Carnforth for the weekend, with a two day working party planned.

As an aside, No 29 was moved out of Deviation Shed to enable the cab roof to be removed. The side tanks are next, as part of the work required to extract the cracked cylinder block which the pattern maker wishes to check. That will involve also the removal of the boiler, for which Chris Cubitt is looking for some strong men (or women) to do the lifting - although he may have to make do with a crane!

The out of date (December 2015) defibrillator pads were also changed.


Saturday 23 January 2016

What a difference the weather makes: no problem this Saturday getting into Grosmont. Another good day getting lots of work done and, more importantly, the Junior Volunteers (JVs) had some two to one training on hand tools with Chris Parrish in the NYMR workshops. Hope the K1 working party had as good a day at Carnforth!

Peter, Brian Armstrong and Andy Lowes went on the NYMR safety briefing to work in the shed yard which they all passed, and have now been invited to do the NYMR PTS course on Sunday 14 February. James Pearcy, Trevor Wilford, Bryan Orange and 4 JVs were working in Deviation Shed while the others were getting educated.

There was though lots of trouble with the pump this time due to rubbish in the pit. Trevor and Peter have cured the leaking joint on the underground sewer pipe, but a non return valve in the bottom of the flexible pipe would be helpful to keep the pump primed.

Progress was made with the proposed new pipes for the axle boxes. James and Andy found the top of the boxes have a ferrule screwed into the top. If the thread is BSP, a nipple can just be screwed in and hopefully the new pipe fitted to this. The right piston was removed before lunch and dimensions recorded. Both piston back faces have been cleaned along with the cylinder studs and faces. The JVs have, when available to work on the locomotive, tapped the cylinder and valve cover nuts and cleaned the washers. The valve packing was removed from the glands and everything cleaned up ready for reassembly.

Bryan Orange, while keeping an eye on the JVs and giving helpful tips and advice, cleaned the valve covers. It was discovered only one has a copper sealing ring fitted: this was annealed by Peter while he was on with the lubrication pipes. The safety valves were removed, being very careful not to pull too hard on the spanner.

In addition, Trevor carried out some burning and welding on No 29 to help out Chris Cubitt.

Lots of bits are now off the engine and ready for the next stage - clean up, decide plan of action and method of repair.

So jobs for Wednesday :-

  • Sweep pit and clear centre drainage channel, also remove stones from pump if necessary;
  • Remove cylinder drain cocks, checking the screw threads and valves;
  • Remove side bar nuts and shims, check clearance with feeler gauge and record on side of rod;
  • Continue with the cab valve maintenance;
  • Boiler doors to clean and mating faces inside the boiler; and
  • Continue frame and wheel cleaning.

Wednesday 20 January 2016

A small but perfectly formed working party on Wednesday - Derek Shorten, Mike Oliver, Jon Bradley and led by Ian Pearson.

Derek continued with cleaning the frames outside and inside when the water in the pit allowed - the pump wasn't working 100%. Jon cleaned all the wash out plugs, down at the MPD, and then made a start on the mudhole/ washout doors. They still require a lot of work and he is planning to come in on Tuesday and Wednesday to continue cleaning the doors. Mike took out both injector steam and stop valves, and also the steam heat valve, and these are now all laid out on the bench in workshop. The valve spindle heads require light machining, and Barry Nesham has agreed to do the machining next week. Ian got back into firebox took out the fusible plugs and cleaned all the grime down from last weeks blow through with compressed air. Subsequently the water hoses were put together, and, with the help of Jon and Derek, the flues were washed out: they are now clear. Finally, the ash was removed from the ashpan and water removed from the smokebox.

Andy will be leading another working party on Saturday to continue the work. Bryan Orange and some of the Junior Volunteers will be attending: could anyone else intending to be present please let Andy know. Remember also that Richard Jackson is planning to hold a safety briefing at 9.30am in the MIC cabin - anyone planning to attend who has not already let Martin Lloyd know, could they please do so as soon as possible.



Page 29 of 30